# Semantic-Aware Adaptive Video Streaming Using Latent Diffusion Models for Wireless Networks

### v1

@@ III-B1 DM-aided Video Streaming
Zhou et al. in [1] presented Codec-aware Diffusion Modeling (CaDM), a novel neural-enhanced video streaming paradigm. CaDM improves compression efficiency by reducing both the resolution and color bit-depth of video frames during encoding. At the decoder side, it employs a denoising diffusion process conditioned on the encoder’s settings to restore high-quality frames, achieving significant bit-rate savings while maintaining superior visual quality. Li et al. in [14] proposed an extreme video compression approach leveraging the predictive power of the forward diffusion process (FDP) with a pre-trained model. However, this method lacks flexibility as it is not adaptable to different video sources. While DMs offer advancements in video streaming, high computational costs and iterative sampling processes can introduce latency, posing limitations for real-time applications.
III-B2 LDM-Aided Video Streaming
@@
Danier et al. in [10] presented LDMVFI, leveraging high-fidelity image synthesis capabilities of DMs for video frame interpolation (VFI). LDMVFI generates dynamic frames based on I-frames and incorporates a vector-quantized auto-encoding model, VQ-FIGAN, to enhance VFI performance. Yu et al. [11] introduced Content-Motion LDM, which decomposes a video into a content frame (image-like) and a low-dimensional motion latent representation. Yu et al. [12] proposed the Projected Latent Video Diffusion Model (PVDM), consisting of an autoencoder stage and a diffusion model stage. The autoencoder employs one latent vector to capture common content (e.g., background) while using two additional vectors to encode motion. By utilizing a 2D image-like latent space, PVDM avoids the computational overhead of traditional 3D convolutional networks, relying instead on a 2D convolution-based diffusion model. Ma et al. in [15] presented DiffVC, a diffusion-based perceptual neural video compression framework that effectively integrates a foundational diffusion model with the video conditional coding paradigm. This framework leverages temporal context from previously decoded frames and the reconstructed latent representation of the current frame to guide LDMs in generating high-quality images. Despite their advancements, these approaches do not consider I-frames and B-frames, relying solely on consecutive frames. This design choice may lead to incorrect I-frame predictions due to improper conditioning, while also reducing training efficiency and increasing bandwidth usage, particularly under complex wireless transmission conditions.
@@
IV Real time Latent Diffusion Adaptive-Bitrate Video Streaming Framework
@@
In this section, we will first propose a channel-aware bitrate selector that automatically decides the bitrate of a video stream for the next chunk according to the channel conditions, semantic performance requirements, and users’ choices in Subsection IV-A. Then, we will introduce the remaining compression, denoising, and reconstruction parts of the latent diffusion adaptive bitrate video streaming (LD-ABS) framework with the selected video resolution to transmit the video frame-by-frame in Subsection  IV-B.
@@
Figure 3: The diagram illustrates the structure of the proposed framework for LDM-enabled real-time video streaming, where $\bm{z}$ is the semantic latent features of video frames $\bm{x}$ encoded by VAE, $T$ represents the total number of forward steps of DM, $t$ denotes the $t$-th step, $n$ represents the $n$-th frame of the video stream with total $N$ frames, $\bm{\theta}$ is the parameters of the DM’s denoising U-Net neural network [4], $r$ represents the starting point for denoising received key frame semantic latent feature signals in the wireless channel, which is related to the channel state, $s$ represents the starting step for fine-tuning non-key frame semantic latent feature, and $\bm{c}$ is the condition for the conditional LDM, i.e., the small-size meta-data being transmitted. Furthermore, $q(\cdot|\cdot)$ denotes the forward process of the DMs, and $p_{\bm{\theta}}(\cdot|\cdot)$ represents the reverse denoising process.
@@ IV-A Channel-Aware Video Streaming Bitrate Selector
  • Neural Network: For each episode $e$, the ABS learning agent determines a suitable bitrate for the next chunk via a neural network (NN). The input features of this NN in $k$-th chunk are ($S_{k}=\{C_{k},M_{k},F_{k}\}$) and are detailed as follows: 1) Past Wireless Channel Estimate: The learning agent uses the past $t$ chunks’ channel status vector $C_{k}=\{c_{k-t},\dots,c_{k}\}$ as input to the NN, where $c_{i}$ represents the channel throughput for video chunk $i$. 2) Latent Content: To detect the diversity of video content, the learning agent uses $M_{k}=\{N_{k+1},V_{k+1}\}$, where $N_{k+1}$ and $V_{k+1}$ denote the size of each bitrate of the next chunk $(k+1)$ and the perceptual quality metrics (semantic ambiguities) for each bitrate of the next chunk, respectively. 3) Player Rollout Playback: Rollout information is collected by the player on the edge user’s side and is defined as $F_{k}=\{v_{k-1},B_{k},D_{k},m_{k}\}$. Here, $v_{k-1}$ represents the video quality of the last video chunk selected, while $B_{k}$, $D_{k}$, and $m_{k}$ represent the buffer utilization, loading time, and the normalized remaining chunks, respectively, for the past $t$ chunks.
@@ IV-B LDM-enabled Compression, Denoising, and Recontruction
For the remainder of LD-ABS, to achieve higher compression of streaming video, we consider using LDM to compress the original reference frames into latent reference I frames. Meanwhile, we only retain the adjustment metadata for the predictive B and P frames. On the user side, latent reference frames are denoised to restore reference frames, and predictive frames are restored using conditional LDM and VFI. Recognizing the challenges faced by LDM-enabled real-time adaptive multi-band streaming, we propose a collaborative end-to-end LDM-enabled streaming framework. This framework aims to deliver high-resolution, low-latency streaming video while conserving communication bandwidth and storage with adaptive time-varying appropriate bitrate. Specifically, the proposed framework involves the following key components:
@@
As depicted in Fig. 3, streaming video is sent to the edge user side via ultra-reliable low-latency communications (URLLC). The fluctuating wireless environment provides an unstable communication bandwidth and channel degradation without prior awareness. ABS must first provide smooth, non-interruptive, low-latency service to satisfy QoS, and then aim to provide high bitrate streaming to satisfy QoE. With a constant channel-bandwidth-ratio (CBR), the transmitter selects the appropriate bitrate based on the wireless environment, as different bitrate streams require distinct latent spaces to store latent frames.
@@
As highlighted in Fig. 3, Step , Upon receiving the L-I-frames $\bm{z}_{r}^{n}$, which contain wireless channel noises and errors, we apply a short reverse process to fine-tune the compressed L-I-frames $\bm{z}_{r}^{n}$ to restore the precise reference frames $\bm{z}_{0}^{n}$. The fine-tuning step aims to improve the reconstruction quality by completing the channel denoising task. Finally, we pass $\bm{z}_{0}^{n}$ through the Red VAE decoder to reconstruct the I-frames. Since the starting point $r$ for channel denoising adapts to the variations in channel estimation results, the proposed LD-ABS approach enhances the robustness of video transmission against challenges such as cahnnel noises, gain attenuation, and semantic errors under time-varying wireless network conditions. Furthermore, since the LDM performs the denoising process in a low-dimensional semantic latent space with fewer steps, it achieves a relatively fast reconstruction speed. This helps mitigate the high computational complexity and slow processing typically associated with DMs.
@@ V Numerical Result and Discussion
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
(e) (f)
Figure 4: Comparison of QoE metrics for different traces: (a)-(b) correspond to FCC traces, (c)-(d) correspond to 3GPP traces, and (e)-(f) correspond to OBOE traces. Specifically, (a),(c),(e) depict trace traffic vs QoE reward, and (b),(d),(f) show the average value of chunk QoE vs CDF.
@@ V-A Experiment setup
To assess performance of average chunk QoE, consistency, and fast adaptation across a wide range of wireless environments, we evaluate our proposed LD-ABS on a virtual player, which is widely assessing system performance of video streaming [6]. This player simulates the adaptive video streaming process using real-world network throughput datasets (3G/HSDPA, FCC, OBOE) and allows comparison with other ABS algorithms under various user and network conditions [6].
@@ V-B Baseline Algorithms
We evaluate LD-ABS against the following state-of-the-art ABS algorithms [6]:
@@ V-C QoE Performance Comparison
Figure 5: Performance evaluation of different ABS algorithms.
@@
The penalty weights are set as $\alpha=1$ and $\beta=2.66$ to balance smooth transitions and minimize rebuffering delays, following [6]. The bitrate selection $a_{k}$ for the $k$-th chunk is optimized to maximize the overall QoE based on the tradeoffs above. The key observation is that LD-ABS consistently outperforms baseline algorithms in terms of average chunk QoE and bitrate utility. Moreover, the results demonstrate that LD-ABS maintains robust performance across all sessions, with the largest proportion of sessions achieving higher QoE values. As shown in Fig. 4(b), which illustrates the cumulative distribution functions (CDFs) of the average QoE across all sessions for different algorithms, at least 95% of LD-ABS sessions achieve an average QoE greater than zero.
Fig. 5 shows that our proposed LD-ABS framework outperforms the testing benchmarks in [6], in terms of chunk QoE, bitrate utility, rebuffering penalty, and smoothness penalty.
Fig. 6 illustrates a sample streaming simulation within a chunk. This chunk consists of one I-frame followed by a P-frame, another P-frame, and a B-frame. Instead of transmitting the entire frames through a noisy wireless channel, only lightweight adjustment metadata is transmitted. The LDM then reconstructs the respective latent frames based on the recovered latent I frame ($z_{o}^{n}$) and the received adjustment metadata, generating the corresponding adjusted frames accordingly.
Figure 6: Sample streaming video chunk simulation on the proposed LD-ABS framework for frame compression, transmission, and reconstruction in wireless networks.
VI Conclusion and Future Direction

### v2

@@ III-B1 DM-aided SemCom for Video Streaming
Zhou et al. in [1] presented Codec-aware Diffusion Modeling (CaDM), a novel neural-enhanced video streaming paradigm. CaDM improves compression efficiency by reducing both the resolution and color bit-depth of video frames during encoding. At the decoder side, it employs a denoising diffusion process conditioned on the encoder’s settings to restore high-quality frames, achieving significant bit-rate savings while maintaining superior visual quality. Li et al. in [11] proposed an extreme video compression approach leveraging the predictive power of the Forward Diffusion Process (FDP) with a pre-trained model.
III-B2 LDM-Aided SemCom for Video Streaming
@@
Danier et al. in [10] presented LDMVFI, leveraging high-fidelity image synthesis capabilities of DMs for Video Frame Interpolation (VFI). LDMVFI generates dynamic frames based on I-frames and incorporates a vector-quantized auto-encoding model, VQ-FIGAN, to enhance VFI performance. Yu et al. [12] introduced Content-Motion LDM, which decomposes a video into a content frame (image-like) and a low-dimensional motion latent representation. Yu et al. [13] proposed the Projected Latent Video Diffusion Model (PVDM), consisting of an autoencoder stage and a DM stage. The autoencoder employs one latent vector to capture common content (e.g., background) while using two additional vectors to encode motion. Ma et al. in [14] presented DiffVC, a diffusion-based perceptual neural video compression framework that effectively integrates a foundational DM with the video conditional coding paradigm. This framework leverages temporal context from previously decoded frames and the reconstructed latent representation of the current frame to guide LDMs in generating high-quality images.
III-C Challenges
Despite the significant potential demonstrated by DMs and LDMs in enhancing SemCom for video streaming, several challenges persist, i.e.,
  • Computational cost and inference latency: Diffusion in pixel space requires dozens to hundreds of denoising steps per frame, resulting in prohibitively high inference latency and processing overhead [11]. Even when the entire diffusion process is moved into a lower-dimensional latent space, full-step denoising remains computationally intensive and cannot consistently meet the low-latency requirements of real-time video streaming.
  • Underutilization of video coding structure (I/B/P-frames): Existing approaches [10, 12, 13, 14] primarily operate on consecutive frames. They often do not explicitly incorporate or leverage the traditional video coding structure involving I-frames and B/P-frames.
  • Poor channel adaptability: Existing works encode and send video at a fixed frame rate and semantic bitrate, without accounting for real-world wireless channel fluctuations. Thus, they cannot maintain stable QoE or temporal consistency when network conditions vary, leading to buffering, frame drops, or inconsistent visual quality.
@@ III-C Challenges
IV Real-Time Latent Diffusion Adaptive-Bitrate Video Streaming Framework
@@
This section first presents a channel-aware bitrate selector that automatically determines the bitrate for the next video chunk based on channel conditions, semantic performance requirements, and user preferences, as described in Subsection IV-A. Next, the compression, denoising, and reconstruction components of the Latent Diffusion Adaptive Bitrate video Streaming (LD-ABS) framework are introduced for frame-by-frame video transmission at the selected resolution in Subsection IV-B. Finally, Subsection IV-C summarizes the motivations and contributions of the proposed framework.
@@ IV-A Channel-Aware Video Streaming Bitrate Selector
  • Neural Network: For each episode $e$, the ABS learning agent determines a suitable bitrate for the next chunk via a NN. The input features of this NN in $k$-th chunk are ($S_{k}=\{C_{k},M_{k},F_{k}\}$) and are detailed as follows: 1) Past Wireless Channel Estimate: The learning agent uses the past $t$ chunks’ channel status vector $C_{k}=\{c_{k-t},\dots,c_{k}\}$ as input to the NN, where $c_{i}$ represents the channel throughput for video chunk $i$. 2) Latent Content: To detect the diversity of video content, the learning agent uses $M_{k}=\{N_{k+1},V_{k+1}\}$, where $N_{k+1}$ and $V_{k+1}$ denote the size of each bitrate of the next chunk $(k+1)$ and the perceptual quality metrics (semantic ambiguities) for each bitrate of the next chunk, respectively. 3) Player Rollout Playback: Rollout information is collected by the player on the edge user’s side and is defined as $F_{k}=\{v_{k-1},B_{k},D_{k},m_{k}\}$. Here, $v_{k-1}$ represents the video quality of the last video chunk selected, while $B_{k}$, $D_{k}$, and $m_{k}$ represent the buffer utilization, loading time, and the normalized remaining chunks, respectively, for the past $t$ chunks.
@@ IV-B LDM-enabled Compression, Denoising, and Recontruction
For the remainder of LD-ABS, to achieve higher compression of streaming video, we consider using LDM to compress the original reference frames into latent reference I frames. Meanwhile, we only retain the adjustment metadata for the predictive B and P frames. On the user side, latent reference frames are denoised to restore reference frames, and predictive frames are restored using conditional LDM and VFI. Recognizing the challenges faced by LDM-enabled real-time adaptive multi-band streaming, we propose a collaborative end-to-end LDM-enabled streaming framework. This framework aims to deliver high-resolution, low-latency streaming video while conserving communication bandwidth and storage with adaptive time-varying appropriate bitrate.
Specifically, the proposed framework involves the following key components:
@@
Figure 3: LDM-enabled adapted bitrate video streaming framework (LDM-ABS)
@@
As depicted in Fig. 3, streaming video is sent to the edge user side via Ultra-Reliable Low-Latency Communications (URLLC). The fluctuating wireless environment provides an unstable communication bandwidth and channel degradation without prior awareness. ABS must first provide smooth, non-interruptive, low-latency service to satisfy QoS, and then aim to provide high bitrate streaming to satisfy QoE. With a constant Channel-Bandwidth-Ratio (CBR), the transmitter selects the appropriate bitrate based on the wireless environment, as different bitrates require distinct latent spaces to store latent frames.
@@
As highlighted in Fig. 3, Step , Upon receiving the L-I-frames $\bm{z}_{r}^{n}$, which contain wireless channel noises and errors, we apply a short reverse process to fine-tune the compressed L-I-frames $\bm{z}_{r}^{n}$ to restore the precise reference frames $\bm{z}_{0}^{n}$. The fine-tuning step aims to improve the reconstruction quality by completing the channel denoising task. Finally, we pass $\bm{z}_{0}^{n}$ through the Red VAE decoder to reconstruct the I-frames. Since the starting point $r$ for channel denoising adapts to the variations of accurately estimated Channel State Information (CSI), the proposed approach enhances the robustness of video transmission against challenges such as cahnnel noises, gain attenuation, and semantic errors under time-varying wireless network conditions. Furthermore, since the LDM performs the denoising process in a low-dimensional semantic latent space with fewer steps, it achieves a relatively fast reconstruction speed. This helps mitigate the high computational complexity and slow processing typically associated with DMs.
@@
IV-C Key Benefits of LD-ABS
Overall, Fig. 3 depicts the structure of the proposed LDM-ABS framework , where $\bm{z}$ is the semantic latent features of video frames $\bm{x}$ encoded by VAE, $T$ represents the total number of forward steps of DM, $t$ denotes the $t$-th step, $n$ represents the $n$-th frame of the video stream with total $N$ frames, $\bm{\theta}$ is the parameters of the DM’s denoising U-Net neural network [4], which include 128 base channels in the base layer of the encoder with channel multipliers [1, 2, 4, 8], $r$ represents the starting point for denoising received key frame semantic latent feature signals in the wireless channel, which is related to the channel state, $s$ represents the starting step for fine-tuning non-key frame semantic latent feature, and $\bm{c}$ is the condition for the conditional LDM, i.e., the small-size meta-data being transmitted. Furthermore, $q(\cdot|\cdot)$ denotes the forward process of the DMs, and $p_{\bm{\theta}}(\cdot|\cdot)$ represents the reverse denoising process.
By combining adaptive bitrate selection with LDM-based compression and reconstruction tailored to different video frame types (I, P, B), LD-ABS provide following key benefits:
  • Enhanced Adaptability: LD-ABS integrates bitrate selector that dynamically adapts streaming to varying network conditions and content resolutions.
  • Reduced Computational Cost and Latency: LD-ABS employs efficient LDMs and avoids lengthy reverse diffusion for every frame, reducing computational load and latency compared to DMs. A short reverse/denoising process is used for received I-frame latents (Fig. 3, Step ), while a conditional LDM fine-tuning process reconstructs B/P-frames based on I-frame latents and motion data.
  • Improved Bandwidth Efficiency and Robustness: Bandwidth usage is reduced by transmitting only compressed I-frame latents (L-I-frames) and compressed motion vectors. Robustness against transmission errors is further strengthened by adaptive I-frame denoising based on channel estimates, while conditional reconstruction of B/P-frames ensures temporal coherence.
@@ V-A Experiment setup
To assess performance of average chunk QoE, consistency, and fast adaptation across a wide range of wireless environments, we evaluate our proposed LD-ABS on a virtual player, which is widely assessing system performance of video streaming [6, 15]. This player simulates the ABS process using real-world network throughput traces (3rd Generation Partnership Project (3GPP), Federal Communications Commission (FCC), OBOE) and allows comparison with other ABS algorithms under various user and network conditions [6]. Edge user-side simulations are performed on a desktop workstation equipped with an Intel Core i7-14700KF CPU, 32GB DDR4 RAM, and an NVIDIA RTX 4070 SUPER GPU (12GB VRAM).
@@ V-B Baseline Algorithms
We evaluate LD-ABS against the following state-of-the-art ABS algorithms [6, 15]:
@@
  • Pensieve: A Deep Reinforcement Learning (DRL)-based algorithm utilizing the A3C method to learn an optimal policy for rate adaptation by mapping buffer occupancy, throughput, and chunk size dynamics.
@@ V-C QoE Performance Comparison
Component 360p 720p 1080p
Transmitter Side
I-frame extraction (①) 1.2 1.8 2.4
VAE encoding (②) 3.4 5.1 6.8
Z-frame compression 1.0 1.2 1.6
Metadata gen. (③) 0.7 1.0 1.3
Receiver Side
Latent decoding (④) 3.6 5.5 6.9
Interpolation (⑤) 2.2 3.6 4.5
I-frame recon. (⑥) 2.8 4.4 5.9
Total Latency 14.9 22.6 29.4
Streaming Method End-to-End Latency
Traditional Broadcasting 6,000 – 13,000
Pixel-Space DDPM 5,000 – 8,000
LD-ABS 2,000 – 3,000
Figure 4: (a) Performance evaluation of different ABS algorithms for FCC traces. (b) Latency breakdown for a chunk (in ms) of LD-ABS at different resolutions. (c) Comparison of end-to-end average total streaming latency (in ms) across different methods.
Figure 5: Cumulative Distribution Function (CDF) of QoE for low-motion and high-motion scenarios across three trace types: FCC, 3GPP, OBOE
@@
The penalty weights are set as $\alpha=1$ and $\beta=2.66$ to balance smooth transitions and minimize rebuffering delays, following [6]. The bitrate selection $a_{k}$ for the $k$-th chunk is optimized to maximize the overall QoE based on the tradeoffs above.
The average CPU and GPU utilization for our edge users is 18% and 41%, respectively—values that are practical and achievable in real-world deployments. Fig. 4(a) shows that our proposed LD-ABS framework outperforms the benchmarks presented in [6] in terms of chunk QoE, bitrate utility, rebuffering penalty, and smoothness penalty. Fig. 4(b) illustrates that the latency for a sample video chunk is well within feasible limits across the processing pipeline shown in Fig. 3. Fig. 4(c) further demonstrates that LD-ABS reduces end-to-end latency compared to traditional approaches.
As depicted in Fig. 5, LD-ABS outperforms the other benchmark with higher average QoE improvements across different traces. The other benchmark frameworks show comparable loss compared with our proposed LD-ABS. With CSI error, LD‑ABS loses <3 % QoE, four times less than the closest competitor. The reason is two‑fold: (i) denoising is performed in a low‑dimensional latent space with only a handful of reverse steps, so the error does not propagate deeply; (ii) the subsequent conditional diffusion for B/P‑frames is anchored to the already‑refined I‑frame, effectively absorbing residual noise. Fig. 6 illustrates a sample streaming simulation within a chunk. This chunk consists of one I-frame followed by a P-frame, another P-frame, and a B-frame. Instead of transmitting the entire frames through a noisy wireless channel, only lightweight adjustment metadata is transmitted. The LDM then reconstructs the respective latent frames based on the recovered latent I frame ($z_{o}^{n}$) and the received adjustment metadata, generating the corresponding adjusted frames accordingly.
Figure 6: Sample video chunk processing using the proposed LDM-ABS framework.
VI Conclusion and Future Directions